A coconut mass 2kg falls from a 30m tall tree. The coconut falls and comes to rest on impact with the person's head who had been sitting directly beneath the tree.

c) The coconut exerted 10,000N of force on the person's head. How long was the coconut in contact?
Answer #1

v = sqrt(2gh) = sqrt(2*9.8*30)

= 24.25 m/s

By impulse-momemtum theorem:

Ft = mv

=> 10,000*t = 2*24.25

=> t = 0.00485 s